# InterMT: Multi-Turn Interleaved Preference Alignment with Human Feedback

### v1

@@ A.5 AI Alignment and RLHF
Aligning LLMs with human preferences is critical for their safe and effective deployment [8]. Among various approaches, supervised fine-tuning (SFT) and reinforcement learning from human feedback (RLHF) have emerged as standard methods for aligning model behavior with human intent [56, 15, 17]. Recent work has extended this alignment framework beyond language-only settings to multimodal scenarios involving both image and text modalities [13, 23, 18, 14]. Such multimodal alignment necessitates addressing challenges like interleaved image-text inputs and outputs, alongside multi-turn interactions that reflect real-world usage. However, the approach to alignment for the multi-turn interleaved multimodal understanding and generation setting still remains an open question.
While highly effective for single-turn instruction following, extending RLHF to multi-turn dialogue introduces significant challenges. These include capturing context-dependent preferences that evolve over the conversation, maintaining long-term coherence and consistency, the increased cost and complexity of collecting high-quality multi-turn preference data, and potential reward hacking where the model optimizes for local turn-level rewards at the expense of overall conversational quality [15].
Some works take an initial step toward multi-turn alignment by leveraging conversation-level human feedback in purely textual multi-turn dialogue, mainly focusing on how to generate better multi-turn dialogue [49]. However, improving instruction-following abilities for multi-modal understanding and generation in multi-turn settings still remains an open challenge.
@@ A.6 Evaluating Multi-turn Multimodal Capabilities
To address this gap, several benchmarks have been proposed [81, 82, 83, 84, 85, 86, 87, 88]. For instance, MMDU introduces a comprehensive benchmark designed to evaluate MLLMs’ abilities in multi-turn and multi-image conversations [82]. It emphasizes the importance of long-context understanding and the integration of multiple images within a single dialogue, pushing models to handle more realistic and complex interactions. ConvBench introduces a hierarchical evaluation framework that assesses LVLMs across three cognitive levels: perception, reasoning, and creativity [89]. This structure enables a nuanced analysis of model performance in multi-turn dialogues, highlighting specific areas for improvement. Similarly, MMMT-IF presents a challenging benchmark focusing on instruction-following in multimodal, multi-turn dialogues [83]. It introduces the Programmatic Instruction Following (PIF) metric, which assesses a model’s ability to follow instructions dispersed across long dialogues, requiring the retrieval and reasoning over instructions spread throughout the context. In the realm of language models, MT-Eval offers a comprehensive benchmark to evaluate multi-turn conversational abilities [90]. By analyzing human-LLM conversations, it categorizes interaction patterns and constructs multi-turn queries to assess models’ performance in maintaining context and coherence over multiple turns. Collectively, these benchmarks highlight the importance of developing evaluation methods that reflect the intricacies of multi-turn, multimodal interactions. However, a common limitation among them is the primary focus on understanding capabilities, often neglecting the generation aspect of multimodal interleaved information. This oversight presents challenges in providing per-turn and overall feedback judgments, which are crucial for the comprehensive assessment and improvement of MLLMs.
@@ Appendix B Data Examples
We conduct an in-depth comparison of both open-source and API-based models, including Janus [91, 92, 93] and Gemini [94, 95], on multi-turn multimodal understanding and generation tasks (Case Study). We further present representative examples of multi-turn QA and preference-annotated instances in InterMT (Examples). Please refer to https://pku-intermt.github.io/ for more details.
@@ C.3 Instituional Review Board (IRB)
We employed 30 full-time crowdsourced workers with substantial experience in multimodal annotation for leading commercial language models. To acknowledge their contributions, we adopted a fair and transparent compensation scheme. The estimated average hourly wage ranged from USD 8.56 to USD 10.23 (XE rate as of 2025/05/13), substantially exceeding the local minimum wage of USD 3.66 in Beijing, PRC [96]. In accordance with local labor laws, workers followed a standard Monday-to-Friday schedule, working eight hours per day with weekends off.
@@ C.4 Comparison with other datasets
CoMM [33] 227k Yes Yes No No 5 No
OBELITICS [72] 141M Yes Yes No No 200 No
MMC4 [69] 101.2M Yes Yes No No 30 No
Visual Dialogue [66] 120k Yes No Yes Yes 80 No
IGC [67] 4.2k Yes No Yes Yes N/A No
Image-Chat [68] 202k Yes No Yes Yes 215 No
MM-Dialogue [65] 1.08M Yes Yes Yes Yes 4184 No
RLHF-V [13] 5.7k Yes No No No - Yes
@@ D.1 Prompt Collection
InterMT is built from 100k image-text QA instances sourced from three primary channels: approximately 72.1% are derived from open-source corpora—namely, publicly available datasets related to vision-language tasks [14, 23, 33] (Table 3 summarizes the open-corpus vision-language datasets used in our pipeline, along with the input-output formats of their original annotations.) ; around 22.8% originate from legally scraped web content (e.g., multimodal platforms such as WikiHow and Pinterest); and the remaining 5.1% are contributed by researcher-curated, human-written prompts. These instances span a wide range of vision-language tasks, including activity generation, data visualization, and table analysis.
@@
LLaVA-Instruct-150K [1]
ART500K [97]
MovieNet [98]
RLHF-V [13]
ShareGPT4V [99]
LLaVA-Instruct-150K [1]
RLHF-V [13]
MM-RLHF [23]
Align-Anything-200K [14]
CoMM [33]
DiffusionDB [100]
MS COCO [101]
HPDv2 [102]
Pick-a-Pic-v2 [61]
Align-Anything-200K [14]
Grounded in theoretical frameworks from linguistics, human-computer interaction, and cognitive psychology [28, 29, 30, 31, 32], we identify five prototypical scenarios that commonly lead to multi-turn conversations in real-world multimodal contexts: (1) incomplete or unclear user cognition; (2) follow-up queries prompted by unsatisfactory initial responses; (3) complex tasks that require incremental, stepwise reasoning; (4) open-ended or companion-like dialogic interactions; and (5) cross-modal mismatches arising from latent inconsistencies between image and text modalities or the need for integrated cross-modal reasoning. Table 4 presents formal definitions of these sceanarios.
@@
Unclear Cognition Based on Belkin’s ASK model, users are in an "anomalous state of knowledge" during retrieval [103]. They recognize knowledge gaps but cannot clearly articulate their needs. Multi-turn dialogue assists in clarifying their goals through guided interaction.
Repeated Attempts Due to Unsatisfactory Answers According to Borlund’s interactive IR model, retrieval is a dynamic, iterative process [104]. Users may re-query after unsatisfactory results. Multi-turn dialogue enables iterative feedback and refinement of information needs.
Complex Tasks Requiring Stepwise Progression Drawing from Ellis’s behavioral model and Kuhlthau’s ISP model, complex tasks require phased progress [105, 106]. Multi-turn dialogue supports task decomposition and information integration, helping users build knowledge step by step.
Exploratory or Companion-like Interaction Bates’s Berrypicking model illustrates non-linear, evolving information behavior [107]. Users follow shifting interests rather than fixed goals. Multi-turn dialogue provides contextual guidance and emotional engagement in open-ended exploration.
@@
Guided by these scenarios, we filter, diversify, and rewrite the original image-text QA instances, resulting in 15,604 unique seed questions, which serve as initial prompts for generating iterative, multi-turn conversations. Figure 11 presents the system prompt used with GPT-4o [38] to evaluate the suitability and potential for multi-turn communication, as well as to assist in filtering and rewriting the original data. Figure 10 illustrates the distribution of seed questions across more than 15 distinct vision-language tasks. Table 5 provides definitions and representative examples for each task category.
@@ D.3 Agent Construction
The agent workflow is built upon a combination of strong open-source models (e.g., Qwen2-VL [35], Qwen2.5-VL [36], Gemma3 [37], and LLaVA-1.5 [4] series) alongside leading API-based models (e.g., GPT-4o [38], Gemini-2.0-Flash [39] and Claude-3.7-Sonnet-Thinking [40]).
Specifically, the following model list are used to construct agents for iterative question generation and response: API-based models include GPT-4o [38], Gemini 2.0 Flash [39], Claude 3.7 Sonnet (both thinking and standard variants) [40]. Open-source models include Qwen-2-VL-72B-Instruct [35], Qwen2.5-VL-32B-Instruct [36], Gemma3-27B-Instruct [37], and LLaVA-v1.5-7B [10].
To support diverse multimodal operations, three types of image-centric tools are integrated: (1) text-to-image generators (e.g., FLUX.1-Schnell [41] and Stable-Diffusion [42]) for producing high-quality images based on prompts; (2) an image editing API (e.g., Gemini-2.0-flash-exp-image-generation [43]) capable of cropping, highlighting, and modifying images; and (3) web-based retrieval interfaces (e.g., Google Images, Pinterest) for sourcing real-world visuals.
@@ D.4 Quality Control and Pruning
  • Image(-Text) Filter: For single-turn image selection, both visual quality and semantic consistency with the text are critical to ensure the selected image is both legible and contextually appropriate. We adopt an image(-text) filter that integrates visual quality assessment and semantic alignment with the input text to rank and filter the candidate image pool returned by the image tool calling module. Specifically, given candidate images $I=\{i_{1},\dots,i_{N}\}$ and caption $T$, we assign each image a score that combines two factors: a rule-based score $\text{Rule}_{j}$ combining multi-dimensions (e.g., resolution, clarity etc.), and a semantic coherence score $\widetilde{\text{Coher}}_{j}$ measuring CLIP-based image–text similarity [108]. The final score is computed as:
@@
  • Consistency Filter: In multi-turn conversations, consistency with prior turns is crucial, both in content (avoiding contradictions with chat history) and style (e.g., maintaining uniform image aesthetics across turns). Advanced models (e.g., GPT-4o [38] and Gemini-2.0-Flash [39]) are employed to better capture such dependencies and ensure coherent filtering across turns.
@@
$j^{*}=\arg\max\nolimits_{j}S(i_{j},T),\quad i_{j^{*}}\text{ as the output (fallback to }i_{1}\text{ if all loads fail).}$).
Thus, each $i_{j}$ is evaluated both on intrinsic visual quality and on semantic alignment with $T$, and the maximum-scoring image is selected.
@@ D.5 Human Preference Annotation
Building on prior works [60, 14, 16], we incorporate human-written natural language feedback into the annotation process. Each feedback instance includes: (i) reason, explaining the rationale behind the assigned score; (ii) critique, identifying strengths and weaknesses of the response based on detailed evaluation criteria; and (iii) refinement, offering suggestions for improving the image or textual quality of the response. The structured feedback protocol significantly improves inter-annotator agreement by 5–10 percentage points, as shown in Table 6.
@@ F.2 Details on Data Labeling Services
Building upon the successes of previous projects such as BeaverTails [58], SafeSora [60], PKU-SafeRLHF [59], and Aligner [16], we again collaborated with the professional annotation service provider AIJet Data. We did not directly interact with the crowdworkers; instead, AIJet managed the entire annotation process. Capitalizing on their expertise in annotating textual data, AIJet curated a dedicated team of experienced annotators tailored to the needs of our project. In light of the task’s complexity, we established a contract with a rate above the industry average to prioritize the engagement of qualified personnel. To ensure consistent annotation quality, we supplied AIJet with comprehensive guidelines aimed at standardizing and refining the labeling criteria.

### v2

@@ A.5 AI Alignment and RLHF
Aligning LLMs with human preferences is critical for their safe and effective deployment [ji2023ai]. Among various approaches, supervised fine-tuning (SFT) and reinforcement learning from human feedback (RLHF) have emerged as standard methods for aligning model behavior with human intent [ouyang2022training, bai2022training, rafailov2024direct]. Recent work has extended this alignment framework beyond language-only settings to multimodal scenarios involving both image and text modalities [yu2024rlhf, zhang2025mm, yu2024rlaif, ji2024align, chen2025ai]. Such multimodal alignment necessitates addressing challenges like interleaved image-text inputs and outputs, alongside multi-turn interactions that reflect real-world usage. However, the approach to alignment for the multi-turn interleaved multimodal understanding and generation setting still remains an open question.
While highly effective for single-turn instruction following, extending RLHF to multi-turn dialogue introduces significant challenges. These include capturing context-dependent preferences that evolve over the conversation, maintaining long-term coherence and consistency, the increased cost and complexity of collecting high-quality multi-turn preference data, and potential reward hacking where the model optimizes for local turn-level rewards at the expense of overall conversational quality [bai2022training].
Some works take an initial step toward multi-turn alignment by leveraging conversation-level human feedback in purely textual multi-turn dialogue, mainly focusing on how to generate better multi-turn dialogue [shani2024multi]. However, improving instruction-following abilities for multi-modal understanding and generation in multi-turn settings still remains an open challenge.
@@ A.6 Evaluating Multi-turn Multimodal Capabilities
To address this gap, several benchmarks have been proposed [xu2024lvlm, liu2024mmdu, epstein2024mmmt, li2024seed, fu2023mme, huang2024dialoggen, hahn2025proactive, zhou2024gate]. For instance, MMDU introduces a comprehensive benchmark designed to evaluate MLLMs’ abilities in multi-turn and multi-image conversations [liu2024mmdu]. It emphasizes the importance of long-context understanding and the integration of multiple images within a single dialogue, pushing models to handle more realistic and complex interactions. ConvBench introduces a hierarchical evaluation framework that assesses LVLMs across three cognitive levels: perception, reasoning, and creativity [liu2024convbench]. This structure enables a nuanced analysis of model performance in multi-turn dialogues, highlighting specific areas for improvement. Similarly, MMMT-IF presents a challenging benchmark focusing on instruction-following in multimodal, multi-turn dialogues [epstein2024mmmt]. It introduces the Programmatic Instruction Following (PIF) metric, which assesses a model’s ability to follow instructions dispersed across long dialogues, requiring the retrieval and reasoning over instructions spread throughout the context. In the realm of language models, MT-Eval offers a comprehensive benchmark to evaluate multi-turn conversational abilities [kwan2024mt]. By analyzing human-LLM conversations, it categorizes interaction patterns and constructs multi-turn queries to assess models’ performance in maintaining context and coherence over multiple turns. Collectively, these benchmarks highlight the importance of developing evaluation methods that reflect the intricacies of multi-turn, multimodal interactions. However, a common limitation among them is the primary focus on understanding capabilities, often neglecting the generation aspect of multimodal interleaved information. This oversight presents challenges in providing per-turn and overall feedback judgments, which are crucial for the comprehensive assessment and improvement of MLLMs.
@@ Appendix B Data Examples
We conduct an in-depth comparison of both open-source and API-based models, including Janus [wu2024janus, chen2025janus, ma2024janusflow] and Gemini [reid2024gemini, team2023gemini], on multi-turn multimodal understanding and generation tasks (Case Study). We further present representative examples of multi-turn QA and preference-annotated instances in InterMT (Examples). Please refer to https://pku-intermt.github.io/ for more details.
@@ C.3 Instituional Review Board (IRB)
We employed 30 full-time crowdsourced workers with substantial experience in multimodal annotation for leading commercial language models. To acknowledge their contributions, we adopted a fair and transparent compensation scheme. The estimated average hourly wage ranged from USD 8.56 to USD 10.23 (XE rate as of 2025/05/13), substantially exceeding the local minimum wage of USD 3.66 in Beijing, PRC [noauthor_undated-nz]. In accordance with local labor laws, workers followed a standard Monday-to-Friday schedule, working eight hours per day with weekends off.
@@ C.4 Comparison with other datasets
CoMM [chen2024comm] 227k Yes Yes No No 5 No
OBELITICS [laurencon2023obelics] 141M Yes Yes No No 200 No
MMC4 [zhu2023multimodal] 101.2M Yes Yes No No 30 No
Visual Dialogue [das2017visual] 120k Yes No Yes Yes 80 No
IGC [mostafazadeh2017image] 4.2k Yes No Yes Yes N/A No
Image-Chat [shuster2018image] 202k Yes No Yes Yes 215 No
MM-Dialogue [feng2022mmdialog] 1.08M Yes Yes Yes Yes 4184 No
RLHF-V [yu2024rlhf] 5.7k Yes No No No - Yes
@@ D.1 Prompt Collection
InterMT is built from 100k image-text QA instances sourced from three primary channels: approximately 72.1% are derived from open-source corpora—namely, publicly available datasets related to vision-language tasks [ji2024align, zhang2025mm, chen2024comm] (Table 3 summarizes the open-corpus vision-language datasets used in our pipeline, along with the input-output formats of their original annotations.) ; around 22.8% originate from legally scraped web content (e.g., multimodal platforms such as WikiHow and Pinterest); and the remaining 5.1% are contributed by researcher-curated, human-written prompts. These instances span a wide range of vision-language tasks, including activity generation, data visualization, and table analysis.
@@
LLaVA-Instruct-150K [liu2024improved]
ART500K [mao2017deepart]
MovieNet [huang2020movienet]
RLHF-V [yu2024rlhf]
ShareGPT4V [chen2023sharegpt4v]
LLaVA-Instruct-150K [liu2024improved]
RLHF-V [yu2024rlhf]
$M$-RLHF [zhang2025mm]
Align-Anything-200K [ji2024align]
CoMM [chen2024comm]
DiffusionDB [wang2022diffusiondb]
MS COCO [chen2015microsoft]
HPDv2 [wu2023human]
Pick-a-Pic-v2 [kirstain2023pick]
Align-Anything-200K [ji2024align]
Grounded in theoretical frameworks from linguistics, human-computer interaction, and cognitive psychology [grice1975logic, grosz1995centering, clark1991grounding, parsing2009speech, traum1995computational], we identify five prototypical scenarios that commonly lead to multi-turn conversations in real-world multimodal contexts: (1) incomplete or unclear user cognition; (2) follow-up queries prompted by unsatisfactory initial responses; (3) complex tasks that require incremental, stepwise reasoning; (4) open-ended or companion-like dialogic interactions; and (5) cross-modal mismatches arising from latent inconsistencies between image and text modalities or the need for integrated cross-modal reasoning. Table 4 presents formal definitions of these sceanarios.
@@
Unclear Cognition Based on Belkin’s ASK model, users are in an "anomalous state of knowledge" during retrieval [belkin1982ask]. They recognize knowledge gaps but cannot clearly articulate their needs. Multi-turn dialogue assists in clarifying their goals through guided interaction.
Repeated Attempts Due to Unsatisfactory Answers According to Borlund’s interactive IR model, retrieval is a dynamic, iterative process [borlund2003concept]. Users may re-query after unsatisfactory results. Multi-turn dialogue enables iterative feedback and refinement of information needs.
Complex Tasks Requiring Stepwise Progression Drawing from Ellis’s behavioral model and Kuhlthau’s ISP model, complex tasks require phased progress [ellis1989behavioral, kuhlthau2004seeking]. Multi-turn dialogue supports task decomposition and information integration, helping users build knowledge step by step.
Exploratory or Companion-like Interaction Bates’s Berrypicking model illustrates non-linear, evolving information behavior [bates1989design]. Users follow shifting interests rather than fixed goals. Multi-turn dialogue provides contextual guidance and emotional engagement in open-ended exploration.
@@
Guided by these scenarios, we filter, diversify, and rewrite the original image-text QA instances, resulting in 15,604 unique seed questions, which serve as initial prompts for generating iterative, multi-turn conversations. Figure 11 presents the system prompt used with GPT-4o [openai2024gpt4o] to evaluate the suitability and potential for multi-turn communication, as well as to assist in filtering and rewriting the original data. Figure 10 illustrates the distribution of seed questions across more than 15 distinct vision-language tasks. Table 5 provides definitions and representative examples for each task category.
@@ D.3 Agent Construction
The agent workflow is built upon a combination of strong open-source models (e.g., Qwen2-VL [Qwen2VL], Qwen2.5-VL [bai2025qwen2], Gemma3 [team2025gemma], and LLaVA-1.5 [liu2024llava] series) alongside leading API-based models (e.g., GPT-4o [openai2024gpt4o], Gemini-2.0-Flash [google2025gemini2flash] and Claude-3.7-Sonnet-Thinking [anthropic2024claude3]).
Specifically, the following model list are used to construct agents for iterative question generation and response: API-based models include GPT-4o [openai2024gpt4o], Gemini 2.0 Flash [google2025gemini2flash], Claude 3.7 Sonnet (both thinking and standard variants) [anthropic2024claude3]. Open-source models include Qwen-2-VL-72B-Instruct [Qwen2VL], Qwen2.5-VL-32B-Instruct [bai2025qwen2], Gemma3-27B-Instruct [team2025gemma], and LLaVA-v1.5-7B [liu2023visual].
To support diverse multimodal operations, three types of image-centric tools are integrated: (1) text-to-image generators (e.g., FLUX.1-Schnell [black2024flux] and Stable-Diffusion [rombach2021highresolution]) for producing high-quality images based on prompts; (2) an image editing API (e.g., Gemini-2.0-flash-exp-image-generation [google2025geminiimagegeneration]) capable of cropping, highlighting, and modifying images; and (3) web-based retrieval interfaces (e.g., Google Images, Pinterest) for sourcing real-world visuals.
@@ D.4 Quality Control and Pruning
  • Image(-Text) Filter: For single-turn image selection, both visual quality and semantic consistency with the text are critical to ensure the selected image is both legible and contextually appropriate. We adopt an image(-text) filter that integrates visual quality assessment and semantic alignment with the input text to rank and filter the candidate image pool returned by the image tool calling module. Specifically, given candidate images $I=\{i_{1},\dots,i_{N}\}$ and caption $T$, we assign each image a score that combines two factors: a rule-based score $\text{Rule}_{j}$ combining multi-dimensions (e.g., resolution, clarity etc.), and a semantic coherence score $\widetilde{\text{Coher}}_{j}$ measuring CLIP-based image–text similarity [radford2021learning]. The final score is computed as:
@@
  • Consistency Filter: In multi-turn conversations, consistency with prior turns is crucial, both in content (avoiding contradictions with chat history) and style (e.g., maintaining uniform image aesthetics across turns). Advanced models (e.g., GPT-4o [openai2024gpt4o] and Gemini-2.0-Flash [google2025gemini2flash]) are employed to better capture such dependencies and ensure coherent filtering across turns.
@@
[…]
Thus, each $i_{j}$ is evaluated both on intrinsic visual quality and on semantic alignment with $T$, and the maximum‐scoring image is selected.
@@ D.5 Human Preference Annotation
Building on prior works [dai2024safesora, ji2024align, ji2024aligner], we incorporate human-written natural language feedback into the annotation process. Each feedback instance includes: (i) reason, explaining the rationale behind the assigned score; (ii) critique, identifying strengths and weaknesses of the response based on detailed evaluation criteria; and (iii) refinement, offering suggestions for improving the image or textual quality of the response. The structured feedback protocol significantly improves inter-annotator agreement by 5–10 percentage points, as shown in Table 6.
D.6 Analysis of Tool Usage Patterns
During the construction of InterMT, we employ a variety of image-related tools, including image generation, editing, and retrieval, to enable backbone models to simulate multimodal, multi-turn input-output scenarios across partial tasks. We analyze the proportion of tool calls across different data categories [^2] Thanks to Reviewer dyyF for inspiring this analysis.. Table 7 summarizes a subset of our results.
Table 7: Tool usage distribution by categories in InterMT.
Category Edit Retrieve Generate
Visual Story Completion 0.12 0.29 0.59
Multi-concept Composition 0.23 0.31 0.46
Content Classification 0.11 0.63 0.26
Visual Analysis 0.34 0.47 0.19
Sequential Image Editing 0.72 0.18 0.10
We observe distinct patterns of tool usage across task categories. For instance, image editing tools dominate sequential image editing tasks, while generation and retrieval tools primarily assist user comprehension. In contrast, all three tools are broadly employed in image tutorial and multi-turn knowledge QA tasks.
@@ F.2 Details on Data Labeling Services
Building upon the successes of previous projects such as BeaverTails [ji2023beavertails], SafeSora [dai2024safesora], PKU-SafeRLHF [ji2024pku], and Aligner [ji2024aligner], we again collaborated with the professional annotation service provider AIJet Data. We did not directly interact with the crowdworkers; instead, AIJet managed the entire annotation process. Capitalizing on their expertise in annotating textual data, AIJet curated a dedicated team of experienced annotators tailored to the needs of our project. In light of the task’s complexity, we established a contract with a rate above the industry average to prioritize the engagement of qualified personnel. To ensure consistent annotation quality, we supplied AIJet with comprehensive guidelines aimed at standardizing and refining the labeling criteria.
F.3 Impact of Image Inputs on Dialogue Dynamics
In InterMT, images simulate realistic user inputs and enhance dialogue understanding and preference annotation in three key ways:
  • Visual Context Improves Queries: Images prompt richer and more complex questions, reflecting real-world interactions.
  • Task-Specific Relevance: In domains such as data visualization or medical imaging, visual signals convey information that text alone cannot.
  • Quality Drop Without Images: Without images, average turns drop from 4.87 to 3.66, and average response length decreases from 874.96 to 402.03 tokens.
To quantify this effect, we conduct an ablation study summarized in Table 8.
Table 8: Impact of image inputs on dialogue dynamics.
Condition Average Turns Average Response Length (tokens)
With Image Input 4.87 874.96
Without Image Input 3.66 402.03
These results highlight the vital role of images in multimodal preference modeling and validate InterMT’s design for realistic evaluation.
F.4 Potential of Extending the InterMT Workflow to Other Modalities
While the current implementation of InterMT focuses on image-text interactions, its underlying design principles naturally extend to a broader range of modalities. In particular, the agent-based workflow and preference modeling architecture are modality-agnostic, providing a flexible foundation for incorporating audio, video, and even haptic or sensor-based inputs. This extensibility stems from the framework’s emphasis on multi-turn, fine-grained human preference modeling, which captures interactional dynamics rather than relying on modality-specific representations.
Future integration of audio and video modalities would enable the framework to more comprehensively reflect the multimodal nature of human communication, capturing not only semantic content but also prosodic, visual, and affective cues that are critical to real-world understanding and alignment. As large foundation models continue to advance toward unified multimodal reasoning, InterMT workflow can serve as a scalable scaffold for evaluating and aligning these capabilities under complex, multi-agent interaction settings. Ultimately, we envision InterMT as a general-purpose paradigm for studying human-AI alignment across modalities, bridging the gap between symbolic dialogue and embodied communication.
F.5 On the Gap Between AI-Synthesized and Human-Human Dialogues
[…]
Consequently, we adopt a strategy that leverages AI for generation and humans for high-level critique and preference annotation, a paradigm consistent with prior work [zheng2023judging, ostyakova2023chatgpt]. This approach allows us to generate a diverse and creative dataset at a feasible cost, a benefit also noted in recent studies [achiam2023gpt, yu2024rlaif]. To mitigate the inherent risks of model-generated content, such as systematic bias and quality degradation, we implement a multi-faceted quality control process.
First, to enhance content diversity and prevent model-specific artifacts, we employ a multi-source generation pipeline. This involves using a mix of closed- and open-source models for language, tool-use, and multimodal understanding, alongside various image generation models (e.g., FLUX, SD-3). We also incorporate a human-in-the-loop diversity audit, where annotators identify and report systematic issues (e.g., repetitive response patterns or image styles), enabling us to refine our models and sampling strategies iteratively.
Second, all generated samples undergo a rigorous two-stage validation process. An initial, automated pre-screening stage filters out candidates using rule-based checks on image metrics (e.g., resolution, clarity) and model-based evaluations like CLIP-based image-text similarity [radford2021learning]. Subsequently, all surviving samples are manually reviewed by our trained annotation team to identify and discard any remaining incoherent or low-quality instances. Only data that successfully pass both automated and human validation are included in the final dataset.